------------------------------------------------------------ Date: Tue Oct 2 08:14:04 2001, Shift: 00:00-08:00 Supervisor: J.I.Jordre, with: Log book#9 pages 74 - 75 Shift Summary: ------------------------------------------------------------ 1. First problems with RHIC "yellow 3.1 cavity", which had to be off during the night. Then yellow sextupole power supply problem, where an expert had to be called for. Therefore no beam until 05:40. Because of the cavity problem, they did not get ZDC rates higher than 160Hz. Now waiting for beam dump and then going to Controlled Access. 2. Run 5086 with 276k trigger 1 events. 3. None ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Tue Oct 02 2001 - 08:14:23 EDT